The perimeter of a rectangle is 76 inches. The length of the rectangle is 5 inches more than twice the width. Find the length and width of the rectangle.
Answer by Liv17(1)

Length (l): 5+2w
Width (w): w
Perimeter formula: 2l+2w
Perimeter = 76
so...
To Find The Width...
76 = 2(2w+5)+2w
76 = 4w+10+2w
76 = 6w+10
66 = 6w
11 = w
To Find The Length Plug In The w Value
length = 5+2w
length = 5+2(11)
length = 27